Effective rate of interest

Find the interest rates earned on each of the following. Round each answer to two decimal places.

  1. You borrow $710 and promise to pay back $781 at the end of 1 year.

    %

  2. You lend $710 and the borrower promises to pay you $781 at the end of 1 year.

    %

  3. You borrow $86,000 and promise to pay back $370,698 at the end of 14 years.

    %

  4. You borrow $20,000 and promise to make payments of $6,687.60 at the end of each year for 5 years.

    %
Answer #1
EAR = ((Ending value / Beginning value)^( 1 / Number of periods)) - 1
a.
EAR =(781/710)^(1/1)-1 10.00%
b.
EAR =(781/710)^(1/1)-1 10.00%
c.
EAR =(370698/86000)^(1/14)-1 11.00%
Year Cash flow
0 -20000
1 6687.6
2 6687.6
3 6687.6
4 6687.6
5 6687.6
=IRR(B292:B297)
EAR 20%
Answers :
a. 10.00%
b. 10.00%
c. 11.00%
d. 20.00%

Answer for a and b are same the only difference is in option a your are borrowing @10% EAR where as in option b you are receiving 10% EAR

10.00% 10.00% 11.00% 284 EAR = ((Ending value / Beginging value)^(1/ Number of periods)) - 1 285 a. 286 EAR =(781/710)^(1/1)-